OpenStudy (anonymous):

Solve for m. -3 + m / 9 = 10 A.-30 B.63 C.87 D.93

OpenStudy (solomonzelman):

\[-3 + m / 9 = 10 \]\[-3+m/9 \color{red} {+3}=10 \color{red} {+3}\]\[m/9=13\]\[(m/9) \color{red} {\times 9}=13 \color{red} {\times 9}\]noww you tell me what m=

OpenStudy (anonymous):

I keep getting the wrong answer